OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r =<8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6> I have NO idea how to work with this question. Help would REALLY be appreciated!

OpenStudy (anonymous):

v=3i-8j-3k w=-4i-2j-6k v.w=(3i-8j-3k).(4i-2j-6k)=(3)(4)+(-8)(-2)+(-3)(-6)=?